Find the slope, y-intercept, and x-intercept of the line that is represented by the equation 3x - 4y= 24 .
Answer by mananth(16949)

3 x + -4 y = 24
Find the slope of this line

-4 y = -3 x + 24
Divide by -4
y = 3/ 4 x -6
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 3/4
yintercept = -6
X intercept , plug y=0
0=(3/4)x-6
3/4 x =6
x=8 , the X intercept